### v2

# The Impact of Train-Test Leakage on Machine Learning-based Android Malware Detection

Abstract
When machine learning is used for Android malware detection, an app needs to be represented in a numerical format for training and testing. We identify a widespread occurrence of distinct Android apps that have identical or nearly identical app representations. In particular, among app samples in the testing dataset, there can be a significant percentage of apps that have an identical or nearly identical representation to an app in the training dataset. This will lead to a data leakage problem that inflates a machine learning model’s performance as measured on the testing dataset. The data leakage not only could lead to overly optimistic perceptions on the machine learning models’ ability to generalize beyond the data on which they are trained, in some cases it could also lead to qualitatively different conclusions being drawn from the research. We present two case studies to illustrate this impact. In the first case study, the data leakage inflated the performance results but did not impact the overall conclusions made by the researchers in a qualitative way. In the second case study, the data leakage problem would have led to qualitatively different conclusions being drawn from the research. We further examine the real-world impact of the data leakage by dissecting the capability of memorization and the capability of generalization of a machine learning model, and show that by removing leakage from testing data, the evaluation results better reflect the machine learning model’s utility in real-world Android malware detection scenarios.
Index Terms:
Android malware detection, Machine learning, Data leakage, Dataset
I Introduction
Machine learning (ML) has been extensively researched for its use in Android malware detection. In such research a raw app is transformed into a suitable numeric representation – a feature vector, a multi-dimensional latent representation, a graph structure, etc., each tailored to a specific ML model. The transformation is typically a lossy function and distinct apps may result in identical app representations. This could lead to a substantial number of app samples in the testing dataset that have representations identical or nearly identical to representations of some samples in the training dataset, a phenomenon known as a type of data leakage [1, 2, 3, 4].
Data leakage has been a recognized phenomenon in ML applications for a long time [5]. Early research defined different types of data leakage [6, 7, 8] which may arise at various stages of ML model development and application. Data leakage has been identified as a significant concern in the implementation of ML algorithms across many scientific fields [1, 9, 2, 3, 4], including in computer security [10]. We focus on the leakage caused by identical or nearly identical app representations between training and testing datasets, a type of leakage that has been largely unexplored in the security research community. In machine learning, a model learned on a training set shall be tested on data that is disjoint from the training data. If an app in the testing set has a representation that is identical to that of an app in the training, the model is tested on data it has already seen during training, and this is called train-test leakage in machine learning literature. Note that such leakage could arise from the inherent nature of the data rather than errors in experimental design. The leakage generally results in inflated performance results that do not reflect the model’s capacity to generalize, since it could simply memorize the leaked data to make a prediction. Some identify the problem as conflating the model’s ability to memorize with its ability to generalize [2]. Even if the representation is not completely identical, but nearly identical, it still calls into question whether the performance measured on the testing data truly reflects the model’s generalizability beyond the data it is trained on. Our framing of the train-test leakage problem follows that used in machine learning literature of domains such as bioinformatics [4], environmental research [9], medical imaging [3], and NLP [2].
Past research on Android malware detection has adopted various approaches to constructing train and test data for machine learning models. Recent work by Pendlebury et al. [11] pointed out biases arising from train/test data construction and how to avoid them. Our analysis found that even when one follows those best practices in constructing training and testing datasets, there could still be large fractions of apps in testing data that share the same feature vectors as apps in training data, leading to data leakage. We ask the questions: 1) how will such leakage impact an ML model’s measured performance, and 2) given that such leakage will happen “naturally” when the model is applied to real-world data, is there any practical value in discerning such differences of performance measurement?
To answer these questions, we present two case studies of recent research efforts of using ML in Android malware detection. The first is that of Chen et al. [12], which investigated the combination of contrastive and active learning in Android malware detection. The second is from the authors of this paper, where we investigate the application of graph neural networks to Android malware detection. Both works divide training and testing data temporally so that training data pre-date testing data. In both cases, we found that substantial fractions of apps in the testing data have identical or nearly identical representations to apps in the corresponding training data. These apps in the testing data are considered leakage based on two different criteria, depending on the specific ML model used.
  1. 1.
    For machine learning models that train on feature vectors extracted from apps, we consider a testing app as leakage if its feature vector is identical to the feature vector of at least one app in the corresponding training data.
  2. 2.
    For machine learning models that analyze app structure through graph representations, these graphs are first converted into numerical latent representations called “graph embeddings” for the final detection stage. To identify data leakage in these models, we leverage cosine similarity, a well-established metric for comparing embeddings. If a testing app’s embedding has a high cosine similarity score with at least one training app’s embedding, we consider the testing app as leakage.
For both Chen et al. [12]’s work and our own work, we evaluate the machine learning models’ performance in two ways: once using the original testing data in the research, and again using the same data but after removing apps identified as data leakage. Our evaluation reveals that removing leakage from the testing data impacts all models from both works. In Chen et al. [12]’s work, all evaluated models experience an average decrease in performance. Similarly, in our work, seven out of eight experiments show a performance decline. In both works, the researchers’ main objectives are to compare various models’ performance. While the models’ average performances dropped when the leakage data was removed from the testing data, we investigated if the comparative results between the models may also be impacted. We found that in Chen et al. [12]’s work, the overall conclusions about the comparative advantage of the proposed contrastive learning+active learning approach still hold, although the extent of improvement did shrink considerably in one of the two datasets. Thus, in this case, the train-test leakage did not impact the overall conclusion from the research in a qualitative way. For our own work on comparing graph neural networks’ capability with a traditional random forest model, the graph neural network model’s performance is inferior to that of the random forest model when measured on the original testing data. After the leakage data is removed from the testing data, the graph neural network’s performance exceeds that of random forest. In this case, the train-test leakage would have led to a qualitatively different conclusion regarding the comparative advantage of the graph neural network model against the random forest model.
In practice a detector needs to make a prediction for every app, regardless of whether the app has been seen before or not. Given this fact, does discerning performance measurement differences after removing the leakage app provide any beneficial insights for operational decision-making? This is especially important in the second case study – does the new measurement suggest favoring graph neural networks over random forests for operational use? How does this reconcile with the fact that random forests performs better than graph neural network on the complete testing data which comprises all apps necessitating processing?
To answer this question, we observe that in practice, machine learning is just one tool out of many analysts rely on to make a determination on an app’s maliciousness. For apps that are known malware, i.e., already labeled as malware in the training data, there is likely other more efficient approach to detect them, e.g., by using a known signature. Only when a malware is unknown, i.e., corresponding to the non-leakage portion of the testing data, a machine learning model is potentially more useful in providing guidance. Thus testing a machine learning model’s performance on the non-leakage portion of the testing data tells more about the model’s utility in practice.
In summary, our research investigates the impact of train-test leakage on the evaluation of machine learning-based Android malware detection systems. In general the measured performance drops when the leakage is removed. Identifying and removing such leakage is not only important in showing machine learning’s true capability of generalizing to substantially different Android apps that it has not been trained on, but also has practical implication on how to apply machine learning in this domain.
II Background
Both traditional machine learning (ML) and deep learning (DL) frameworks have been proposed and extensively investigated for Android malware detection and demonstrated remarkable performance over the past decade. We provide some background relevant to the discussion in this paper.
II-A App Representation for Machine Learning
[…]
DL-based Android malware detection has also used vectorized app representations as model inputs [13, 14, 15, 16]. However, DL’s input capability extends beyond one-dimensional, allowing for richer multi-dimensional tensor representations of text, image, graph, and other structured data as inputs. Furthermore, DL has the ability to automatically learn latent vector representations (called embeddings) that capture high-level semantic relationships in its inputs [17, 18]. Prior works have transformed raw bytecode of an APK’s DEX files into a grayscale or color image suitable for convolutional neural networks, which subsequently learn feature representations at many levels [19, 20]. Other works have represented an app as method-level Control Flow Graphs (CFGs) or Data Flow Graphs (DFGs), which can then be fed into convolutional neural networks or graph convolutional network architectures [21, 22, 23].
II-B Experimental Biases
Despite significant advances in machine learning based Android malware detection, experimental biases and pitfalls persist, demanding careful consideration. Pendlebury et al. [11] highlighted two crucial concerns: 1) spatial bias: test data does not reflect real-world ratio of benign to malicious apps, and 2) temporal bias: apps in training and testing do not follow a realistic timeline. One example of temporal bias is when some apps in the training set are more recent than those in the testing set, essentially using future apps to predict the behavior of earlier apps. This bias often occurs when a large amount of data is collected from single or multiple sources and randomly split into training and testing sets without considering the release date of the apps. Models evaluated on such temporally biased data cannot reflect its utility in operation, where such future knowledge is impossible to obtain. Furthermore, Arp et al. [10] identified sampling bias as another pitfall in learning-based security systems, including Android malware detection. This bias occurs when sampling inadvertently introduces artifacts that do not exist in real-world data. For example, if benign apps are primarily collected from one market, while malicious apps are mainly collected from others, the model may end up learning the difference between the two markets instead of the difference between malware and benign apps. Dataset collection needs to be carefully done to avoid these biases.
III Train-Test Leakage in Android Malware Detection
For machine learning models, if the representation of an Android app in the testing set coincides with the representation of another app in the training set, the app in the testing set is considered leakage – the model will be tested on an app that the model was also trained on. Even though the underlying apps are distinct, since an ML model only learns from the app’s representation, identical representations are the same data to the ML model. Train-test leakage has been extensively discussed for machine learning models applied in other domains, including bioinformatics [4], environmental research [9], medical imaging [3], and NLP [2]. It is well known that such data leakage will inflate performance results, especially if the fraction of leakage is significant. The main problem is that performance metrics measured on the testing set containing leaked data conflates the model’s ability to generalize with its ability to memorize. For the portion of apps in the testing data that are identical to some in training data, the model can simply memorize those data’s label as “prediction” result. This could undermine the validity of conclusions drawn from ML-based research, especially if the conclusion is sensitive to a model’s ability to generalize to data it has not seen in training. In these domains, researchers have addressed the train-test leakage by re-dividing the train/test splits to avoid such leakage. This is not an option for most security problems, and for Android malware detection in particular, since doing so will likely introduce temporal biases as discussed in Section II-B. To assess the impact of data leakage on evaluation results, we can exclude the “leaked apps” from the testing dataset, re-evaluate the models on the “cleaned” testing data, re-draw all conclusions thereafter, and compare with the results before the leakage removal. To the best of our knowledge, there has not been any study that directly examines train-test leakage’s impact on evaluating machine learning models’ utility in a security problem domain. One recent manuscript [24] identifies train-test leakage in vulnerability datasets for evaluating code language models, but did not isolate the leakage’s impact on evaluating the models’ performance.
Another type of data leakage, which is sometimes confused with train-test leakage, is called temporal leakage [1]. Temporal leakage arises from temporal bias as discussed in Section II-B. It happens when testing data pre-dates training data, essentially requiring a model to be trained on future data not available at the time of model usage. Temporal leakage is always a flaw in experiment design, since it is evaluating a model under a scenario that can never materialize in real world. Train-test leakage, on the other hand, could very well happen in a realistic model application scenario. For example, due to the nature of Android app evolvement, apps released today may bear similarity to some apps released in the past, creating possible train-test leakage when those older apps are used for training. The problem caused by train-test leakage is the conflation of model’s ability to generalize with its ability to memorize, leading to potentially inaccurate conclusion drawn about a machine learning model’s utility in practice. This paper’s sole focus is train-test leakage. The two case studies we use are both free of temporal leakage.
III-A Evaluation Metrics
F1 score is a common metric adopted by Android malware detetion work, but it is sensitive to the malware/benign app ratio [11]. When we remove leakage apps from testing set to examine the leakage apps’ impact on the model’s performance measurement, we may also have changed the malware ratio in the testing set, which may impact F1 score by itself. For isolating the impact of train-test leakage on the performance measurement, we use balanced accuracy (BA), defined as the arithmetic mean of sensitivity and specificity. The full definitions of the various evaluation metrics can be found in Appendix A. Unlike F1 score, BA is insensitive to class ratio, which will allow us to show the impact of the leakage on the performance measurement. We can compare an ML model’s BA on the testing data before and after the leakage is removed. A performance decrease after removing the leakage indicates that the leaked apps in the test data inflated the performance results of the model.
In the next two sections, we present two case studies of research where ML is used for Android Malware detection. We examine the impact of leaked apps in testing data on the performance results, and discuss how this may affect the ultimate conclusions the research aims to draw. When comparing the performance results before and after leakage removal, we include both balanced accuracy (BA) and the original metrics adopted by the researchers (e.g., F1 score).
IV Case Study 1: Continuous Learning for Android Malware Detection
Chen et al.’s work [12] is one of the most recent studies that apply machine learning to Android malware detection. The authors made all research artifacts publicly available, including the datasets used in the experiments and the model-related code. We were able to reproduce the results published in the paper using the published code and data without any trouble. This enabled us to examine the effect of train-test data leakage on this work.
IV-A Summary of the Research and Datasets Used
Chen et al. [12] proposed a continuous learning approach that combines contrastive learning and active learning to improve ML-based Android malware detector’s capability of adapting to concept drift [25]. To address the challenge posed by evolving malware techniques, the researchers introduced a hierarchical contrastive learning scheme and a novel sample selection technique for continuous learning as shown in Figure 1. The process begins by using the apps from the first year of the dataset as the initial training data to build the initial hierarchical contrastive classifier. Then an active learning process starts. The trained classifier is used to predict the labels of next month’s samples. The classifier computes an uncertainty score for all samples in the month (based on contrastive learning embeddings) and selects the apps with the most uncertain scores, based on a labeling budget. These selected apps are then labeled and added into the training set for the next iteration, where a new model is trained using warm start (i.e., starting with the model from the prior iteration). This process repeats and the model is continuously trained and tested on each next month’s apps. The ultimate goal is to minimize human labeling effort, while keeping the model updated with evolving malware.
Figure 1: The left rectangle illustrates Chen et al.’s [12] continuous learning approach. Each month new samples are added to the training data through active learning; a new hierarchical contrastive classifier is trained using warm start and evaluated on next month’s data. Month 13-18 are used as validation data. Test data starts from month 19. The dotted rectangle on the right illustrates removing leakage apps from test data for evalaution. The non-leak evaluation does not change Chen et al.’s original approach; it simply removes leakage apps from testing data for evaluation.
The hierarchical contrastive classifier proposed by the authors consists of two subnetworks. The first subnetwork is an encoder (Enc), which encodes the original inputs using contrastive learning based on malware family information. The second subnetwork is a multilayer perceptron (MLP), which uses the encoded input to predict a binary label. This model architecture is denoted as Enc+MLP. The authors compared their approach with several active learning baselines using both binary and multi-class classifiers. The binary classifiers include a fully connected neural network (NN), a linear support vector machine (SVM), and gradient boosted decision trees (GBDT), and the multi-class classifiers (which predict malware families) [^2] The authors note that the output of the multi-class classifiers is converted to binary classes (malware/benign) to allow for a fair evaluation of all approaches used. include multiplayer perceptron (MLP) and SVM. An additional experiment investigated a “Multi-class MLP + Binary SVM” architecture.
Chen et al. [12] compared their continuously learning approach against these prior classifiers combined with several active learning schemes. Two datasets were used in the evaluation: APIGraph [26], with app years 2012-2018, and AndroZoo [27], with app years 2019-2021. The apps are represented as feature vectors, using DREBIN [28] features extracted from the apps, which capture an app’s access to hardware, requested permissions, component names, intents, restricted API calls, used permissions, suspicious API calls, and network addresses. The number of DREBIN features used in the APIGraph dataset is 1,159, while the number of features used in the AndroZoo dataset is 16,978 [12]. For the APIGraph dataset, the initial model was trained on year 2012 data, and validation was performed on data covering the months from 2013-01 to 2013-06. The testing data included the months from 2013-07 to 2018-12. For the AndroZoo dataset, the initial model was trained on year 2019 data, while validation was performed on the months from 2020-01 to 2020-06, and testing included the months from 2020-07 to 2021-12. This arrangement of training/validation/testing data ensured that no temporal leakage was introduced. The authors used the metrics of FNR, FPR, and F1 score. In both datasets, the performance metrics were averaged across all testing data months. The results show that their approach outperforms all prior works.
(a) APIGraph test data (2013-07 to 2018-12)
(b) AndroZoo test data (2020-07 to 2021-12)
Figure 2: Leakage ratio in APIGraph and AndroZoo datasets used in Chen et al.’s work
IV-B Identification of Train-Test Leakage
In identifying train-test leakage in the dataset, we consider a test app a leakage if its feature vector is identical to that of at least one app in the corresponding training data. In the continuous learning approach, new samples are added to the training data each month, and the next month’s data is used for testing. Thus, to identify leakage in a testing month, we first need to track all the newly labeled samples added to the training data in the previous months. We then compare the apps in the testing data against all the apps in the initial training data (first 12 months) plus the newly labeled apps that have been added so far. Any app in the testing month that share the same feature vector with at least one of these training apps is identified as leakage.
Chen et al. [12] evaluated seven different models, combined with several active learning schemes for selecting uncertain apps for labeling and adding into training data. Thus, for the same testing month, the training data used depends on which model is being evaluated and what sample selection scheme is used in active learning. All models use the first 12 months of data as initial training, but the newly added samples for subsequent training will differ across the different models. In effect, for the same testing month, each model being evaluated gives rise to a distinct subset that is considered leakage. To examine the impact on a model’s performance, we would need to remove the model-dependent leakage apps. This poses a problem for comparing the performance across the models: the testing data is no longer the same for those models after the leakage is removed, since what is removed may differ from one model to another. This will make it impossible to compare the performance results of those models, since such comparison ought to be done on the same testing data.
To allow us to compare the various models after removing leakage, we must adhere to two principles:
  1. 1.
    All leakage apps in the testing month must be removed for all models.
  2. 2.
    All models must be evaluated on the same testing data.
It then follows that we must remove from the testing data the union of the leakage apps for all models, to allow us to compare the results among those models after leakage removal. The dotted rectangle at the right corner of Figure 1 illustrates the removal of the union of the leakage apps for all models from a testing month. We calculate the ratio of the removed apps over the total number of apps in that test month. Figure 2(a) and Figure 2(b) illustrate this leakage ratio for each testing month. Higher monthly labeling budget leads to more apps added to the training data over time, and hence higher chance a testing app’s feature vector being identical to one of the training apps. Overall about 10%-40% of testing apps are found to be leakage for the APIGraph test data, and about 10%-30% of testing apps are found to be leakage for the AndroZoo test data.
It is important to note that the removal of leakage apps is only done for the purpose of evaluating the trained models. We do not change how the active learning scheme works in Chen et al.’s work. For each month, the total set of apps in that month is used for identifying the most uncertain predictions, just as what the authors did in their experiment. We simply provide another measurement of the models’ performance – running the same models on the testing data with leakage removed, and compare this performance against the performance measured on the unchanged testing data.
IV-C Impact of the Leakage on Performance Results
TABLE I: Results Before and After Removing Leak from Test Data in Chen et al.’s Work – APIGraph Dataset
Monthly
Sample
Budget
Model
Architecture
Measurement on Original Data
Average Performance (%)
Measurement after Leak Removed from Test
Average Performance (%)
FNR FPR F1 BA FNR FPR F1 BA
50 Binary MLP 23.99 0.56 83.37 87.72 27.44 0.54 80.47 86.01
Multiclass MLP 14.57 5.58 72.57 89.92 18.27 4.33 69.38 88.70
Multiclass MLP
+ Binary SVM
38.87 1.40 69.49 79.87 38.74 1.24 68.37 80.01
Binary SVM 16.92 0.61 87.72 91.23 20.78 0.63 84.24 89.30
Multiclass SVM 27.94 0.17 82.64 85.94 35.72 0.18 76.23 82.05
Binary GBDT 31.75 0.54 77.92 83.85 30.66 0.76 76.61 84.29
Enc + MLP 14.18 0.64 89.39 92.59 19.19 0.50 86.08 90.15
($\downarrow$ 2.74) ($\uparrow$ 0.03) ($\uparrow$ 1.67) ($\uparrow$ 1.36) ($\downarrow$ 1.59) ($\downarrow$ 0.13) ($\uparrow$ 1.84) ($\uparrow$ 0.85)
100 Binary MLP 20.65 0.46 86.10 89.44 23.91 0.44 83.24 87.82
Multiclass MLP 14.44 4.43 75.42 90.57 15.81 4.49 70.34 89.85
Multiclass MLP
+ Binary SVM
28.37 1.59 76.38 85.02 30.45 1.57 72.10 83.99
Binary SVM 15.06 0.68 88.56 92.13 17.91 0.68 85.49 90.71
Multiclass SVM 28.36 0.17 82.18 85.73 28.36 0.18 81.90 85.73
Binary GBDT 27.76 0.67 80.15 85.79 26.62 0.77 79.20 86.30
Enc + MLP 12.43 0.53 90.78 93.52 16.20 0.45 88.18 91.68
($\downarrow$ 2.63) ($\downarrow$ 0.15) ($\uparrow$ 2.22) ($\uparrow$ 1.39) ($\downarrow$ 1.71) ($\downarrow$ 0.23) ($\uparrow$ 2.69) ($\uparrow$ 0.97)
200 Binary MLP 18.96 0.47 87.17 90.29 22.31 0.47 84.28 88.61
Multiclass MLP 13.03 4.69 75.45 91.14 16.51 4.00 71.70 89.74
Multiclass MLP
+ Binary SVM
29.93 1.75 74.48 84.16 34.50 1.37 70.38 82.07
Binary SVM 13.76 0.86 88.64 92.69 16.55 0.93 85.16 91.26
Multiclass SVM 21.19 0.21 86.90 89.30 24.85 0.25 83.84 87.45
Binary GBDT 24.71 0.56 82.71 87.37 25.24 0.69 80.83 87.03
Enc + MLP 10.18 0.49 92.38 94.66 14.37 0.56 88.70 92.54
($\downarrow$ 3.58) ($\downarrow$ 0.37) ($\uparrow$ 3.74) ($\uparrow$ 1.97) ($\downarrow$ 2.18) ($\downarrow$ 0.37) ($\uparrow$ 3.54) ($\uparrow$ 1.28)
400 Binary MLP 14.49 0.62 89.22 92.44 20.63 0.47 85.40 89.45
Multiclass MLP 14.87 3.45 77.98 90.84 17.58 3.37 73.55 89.53
Multiclass MLP
+ Binary SVM
24.48 1.88 77.86 86.82 28.20 2.01 72.22 84.90
Binary SVM 12.86 0.90 89.02 93.12 15.69 1.06 85.35 91.63
Multiclass SVM 17.87 0.24 88.88 90.95 22.75 0.30 84.87 88.47
Binary GBDT 20.16 0.46 86.24 89.69 21.91 0.66 83.40 88.71
Enc + MLP 8.67 0.44 93.33 95.44 14.21 0.59 88.67 92.60
($\downarrow$ 5.82) ($\downarrow$ 0.18) ($\uparrow$ 4.11) ($\uparrow$ 3.00) ($\downarrow$ 6.42) ($\uparrow$ 0.12) ($\uparrow$ 3.27) ($\uparrow$ 3.15)
TABLE II: Results Before and After Removing Leak from Test Data in Chen et al.’s Work – AndroZoo Dataset
Monthly
Sample
Budget
Model
Architecture
Measurement on Original Data
Average Performance (%)
Measurement after Leak Removed from Test
Average Performance (%)
FNR FPR F1 BA FNR FPR F1 BA
50 Binary MLP 51.55 0.30 60.92 74.08 58.82 0.40 53.40 70.39
Multiclass MLP 54.84 15.29 33.45 64.94 37.82 32.78 31.09 64.70
Multiclass MLP
+ Binary SVM
72.55 2.53 34.80 62.46 83.58 1.49 22.92 57.46
Binary SVM 48.77 0.29 63.42 75.47 57.56 0.28 54.28 71.08
Multiclass SVM 64.96 0.12 47.89 67.46 72.50 0.09 39.16 63.70
Binary GBDT 50.35 0.47 61.06 74.59 58.34 0.48 52.48 70.59
Enc + MLP 29.79 0.53 79.29 84.84 55.65 0.47 56.83 71.94
($\downarrow$ 18.98) ($\uparrow$ 0.24) ($\uparrow$ 15.87) ($\uparrow$ 9.37) ($\downarrow$ 1.91) ($\uparrow$ 0.19) ($\uparrow$ 2.55) ($\uparrow$ 0.86)
100 Binary MLP 45.79 0.31 65.80 76.95 58.22 0.29 54.08 70.75
Multiclass MLP 42.23 21.24 36.59 68.27 56.76 20.54 27.61 61.35
Multiclass MLP
+ Binary SVM
77.32 1.59 31.51 60.55 80.44 2.01 25.78 58.77
Binary SVM 43.07 0.32 68.33 78.31 52.28 0.29 59.23 73.72
Multiclass SVM 53.97 0.06 59.14 72.99 64.04 0.10 47.90 67.93
Binary GBDT 48.59 0.76 62.58 75.33 57.56 0.65 53.62 70.90
Enc + MLP 30.17 0.44 79.46 84.70 47.19 0.37 64.94 76.22
($\downarrow$ 12.90) ($\uparrow$ 0.12) ($\uparrow$ 11.13) ($\uparrow$ 6.39) ($\downarrow$ 5.09) ($\uparrow$ 0.08) ($\uparrow$ 5.71) ($\uparrow$ 2.50)
200 Binary MLP 43.58 0.27 67.63 78.07 56.16 0.22 56.12 71.81
Multiclass MLP 32.42 24.78 37.74 71.40 46.85 35.64 26.90 58.76
Multiclass MLP
+ Binary SVM
65.71 1.75 42.32 66.27 75.32 0.90 33.92 61.89
Binary SVM 40.31 0.37 70.24 79.66 48.94 0.37 61.80 75.34
Multiclass SVM 46.55 0.16 65.56 76.64 59.44 0.14 52.11 70.21
Binary GBDT 42.97 0.80 67.28 78.12 53.07 0.72 58.28 73.11
Enc + MLP 29.40 0.34 79.35 85.13 43.21 0.43 68.08 78.18
($\downarrow$ 10.91) ($\uparrow$ 0.03) ($\uparrow$ 9.11) ($\uparrow$ 5.47) ($\downarrow$ 5.73) ($\uparrow$ 0.06) ($\uparrow$ 6.28) ($\uparrow$ 2.84)
400 Binary MLP 35.96 0.41 73.99 81.81 42.33 0.24 69.55 78.72
Multiclass MLP 24.98 29.96 37.56 72.53 36.28 38.09 30.23 62.82
Multiclass MLP
+ Binary SVM
61.55 1.67 46.81 68.39 81.04 1.29 27.14 58.84
Binary SVM 34.73 0.43 74.12 82.42 43.18 0.46 66.39 78.18
Multiclass SVM 41.75 0.13 69.59 79.06 51.80 0.08 59.54 74.06
Binary GBDT 33.62 0.38 76.82 83.00 42.95 0.39 69.12 78.33
Enc + MLP 23.59 0.37 84.23 88.02 35.51 0.52 73.73 81.98
($\downarrow$ 10.03) ($\downarrow$ 0.01) ($\uparrow$ 7.41) ($\uparrow$ 5.02) ($\downarrow$ 6.82) ($\uparrow$ 0.28) ($\uparrow$ 4.18) ($\uparrow$ 3.26)
We examined how the presence of train-test leakage might impact the main research conclusion drawn by the authors. We did so by evaluating the models both on the original test data, and on the test data after removing the leakage. We then compared the results. We expected that all models’ performances would drop after leakage in test data is removed. The key question is whether the main conclusion in the paper, that the proposed continuous learning approach outperforms the prior approaches, will still hold after leakage is removed from test data. Before we conducted the experiments, we first ran the model code on the datasets published by the authors to see whether we could replicate similar performance results as reported in Chen et al.’s original paper [12]. Apart from small differences in the performance numbers due to differences in numerical precision across various hardware architectures and non-deterministic operations related to parallelism and threading333 Due to these factors the model might not produce exactly the same results even when run on the same data with the same settings twice., our results replicated those reported in the original paper.
We then re-ran all the experiments to obtain the new performance measurement on testing data with leakage removed. We kept the rest of the experiments and setup the same as in the original paper [12]. Tables I and II show the performance results of the various models before and after the train-test leakage is removed. We follow the same table format as that used in the original paper [12], and add a set of columns for performance metrics measured on the test data when the leakage has been removed. The original paper used FNR, FPR, and F1 score. We also included Balanced Accuracy (BA) as an additional metric. In both tables, all columns under “Measurement on Original Data” are produced by our own experiments replicating the results from the original paper. All columns under “Measurement after Leak Removed from Test” are results obtained after we remove leakage from the testing apps before measuring the models’ performance. The table is organized by blocks for monthly sample budget from 50 to 400. For a fixed monthly sample budget, the authors’ approach (Enc + MLP) is compared with the best performance from the prior approaches, both indicated in bold font. The value in the colored parentheses with up or down arrows indicate the difference between the Enc + MLP approach and the best prior approaches, with green in favor of Enc + MLP and red in favor of the best prior approach.
For both APIGraph and AndroZoo datasets, there is a consistent decrease of the overall model performance when leakage is removed from test data. The F1 metrics overall show decrease after leakage is removed from testing data. But F1 metric is also sensitive to the class ratio in the testing data. If the leakage removal reduces the malware ratio in the testing data, this would result in a decreased F1 score even if the model’s predicting capability remains the same. As discussed in Section III-A, the BA metric isolates the performance results from the change in the malware ratio in the test set. By comparing the BA numbers before and after leakage removal, one can see how much performance is inflated by the train-test leakage. We can see that the two BA metric columns for the same row almost always show a decrease when leak is removed from test. This decrease is more notable for the AndroZoo data (Table II) than for the APIGraph data (Table I). This confirms our hypothesis that train-test leakage inflates a machine learning model’s measured performance.
We now examine whether the inflated performance results due to train-test leakage would impact the overall research conclusion, in this case whether Enc+MLP outperforms the prior approaches. We notice that with only one exception (FPR number of monthly labeling budget 400 in both tables) the color-coded comparison results remain the same color before and after leakage removal. If we only look at the comparative results for the overall metrics of F1 and BA, the color-coded differences all show the same green color, meaning the Enc+MLP outperforms the best from the prior baseline approaches. Thus, while the train-test leakage inflated the performance results of all models, the relative performance between Enc+MLP and the prior models, as measured by F1 score and BA, remain the same. The overall conclusion from Chen et al.’s work still holds. We do notice, however, that for the AndroZoo data (Table II), the Enc+MLP approach’s improvement margin from best prior works considerably shrinks after leakage is removed from test data.
V Case Study 2: Graph Convolutional Networks for Android Malware Detection
Our second case study is research of the authors of this paper. We have investigated the use of graph convolutional networks (GCN) for Android malware detection. We apply GCN on Android apps’ interprocedural control flow graphs (ICFG) to build a deep learning model for Android malware detection, and compare the performance of the GCN model against that of random forest, a commonly used traditional ML model. In this section, we report on the train-test leakage identified in this research of our own, and how it impacts the conclusion we could draw from the research.
V-A Summary of the Research and Dataset Used
(a) Breakdown of a sliding window
(b) Data window sliding
Figure 3: Data Sliding Windows
GCNs leverage graph structures as input, which allows for potentially capturing richer static code semantics, such as that expressed in an Android app’s ICFG. We use an Android app’s ICFG444 To obtain a single ICFG for an Android app, we combine the ICFG of each component of the app based on the life cycle semantics of how the methods of various components are called by the Android runtime. as the GCN model’s input, where each node represents an API call, with additional attributes including class path, method name, parameter type, and node degree. Following the standard GCN training process, we obtain a graph embedding for the ICFG, achieved through multiple rounds of message passing within the GCN, where each node’s embedding is updated based on its neighbors. Finally, a readout layer aggregates these node embeddings into a unified graph embedding for the entire ICFG. This component of the model can be seen as a graph encoder. The unified graph embedding learned by the encoder is then fed as input to the classification layer of the model. More details about the GCN model can be found in the Appendix B.
We compare the GCN approach to a standard ML approach – random forest (RF). The RF model takes as input a feature vector representing an app. We extract DREBIN-like static features and rank them in decreasing order of their mutual information with the class label. We then select 2,970 most informative features for training an RF model, which include permissions, intent actions, discriminative APIs, obfuscation signatures, and native code signatures. The mutual information is close to 0 for features ranked lower.
[…]
The complete dataset encompasses 24 batches, totaling 5,760 malicious and 90,240 benign apps, all chronologically ordered by their release date on Google Play. This dataset provides eight sliding windows (the window slides two batches each time) for evaluating machine learning models through different training, validation, and testing data. This sliding window approach, as the sampling approach for batch construction, also avoids the temporal bias [11] as discussed in Section II-B.
V-B Identification of Train-Test Leakage
Figure 4: Ratio of leakage apps in sliding-window-based test data (2018-01 to 2021-12)
The two machine learning models our research compares, GCN and RF, use different input representations for an Android app. In GCN, an input app is represented as an ICFG adjacency matrix, together with a list of attributes for each node, whereas in RF an input app is represented as a feature vector. In prior discussion, we have explained the definition of train-test leakage when a feature vector is used as an app’s representation. These are apps in the testing data that share identical feature vectors with at least one app in the training data. We denote these leakage apps as $Leak_{\textit{fv}}$. While this definition works well for RF, it is not suitable to use in identifying train-test leakage for the GCN-based models since there an app is represented as a graph. Given that there is currently no known efficient algorithm to solve graph isomorphism, it is not clear how to efficiently determine if two Android apps’ ICFGs are identical. However, as mentioned above, the GCN model consists of an encoder and a classification layer. The encoder generates a graph embedding, i.e., a vector of floating numbers, which is provided as input to the classification layer.
Instead of comparing the ICFGs directly, we have chosen to use the computed graph embeddings of the ICFGs to identify train-test leakage. Specifically, given the embeddings $emb_{1}$ and $emb_{2}$ of two Android apps’ ICFGs, we compute cosine similarity between the two embeddings as
$\textit{cosine similarity}=\frac{emb_{1}\cdot emb_{2}}{\|emb_{1}\|\|emb_{2}\|}$
where $\cdot$ is the inner product operation and $\|v\|$ represents the length of a vector $v$. The cosine similarity measures how similar the two vectors are and ranges from 0 to 111, where 111 means the two vectors are identical, while 0 means that the two vectors are orthogonal and share no similarity.
Note that due to GCN model’s stocasticity and the numerical precision limitations, even if we feed the exact same ICFG multiple times to a GCN, the model will not generally produce the exact same graph embeddings. Thus, for the purpose of identifying train-test leakage, we chose a threshold $M$. A test app is considered a leakage if there is at least one training app whose cosine similarity with the testing app is at least $M$. We denote these apps as $Leak_{\textit{emb}}$. For a given testing set, we choose $M$ so that the two sets of leakage apps in the testing data, $Leak_{\textit{fv}}$ and $Leak_{\textit{emb}}$, are most aligned. The $M$ value varies from one testing window (Figure 3) to another, but they are all very close to 111 (e.g., 0.99980.99980.9998). More details on how the threshold $M$ is determined can be found in Appendix C.
Figure 4 illustrates the ratio of leakage apps ($Leak_{\textit{fv}}$, $Leak_{\textit{emb}}$, and their union) on the testing data used in our experiments. A significant portion of apps in the testing data (oftentimes exceeding half of the total apps) are identified as train-test leakage. Since the two models give rise to two different sets of leakage apps, we need to remove the union of the two leakage sets from the testing data in order to compare the two models’ performance after leakage removal, for the same reasons as explained in Section IV-B.
V-C Impact of the Leakage on Performance Results
TABLE III: Compare GCN and RF Before and After Leakage Removal
Window
Model
Architecture
Measurement on Complete Test Data
Performance (%)
Measurement after Leak Removed from Test
Performance (%)
FNP FPR F1 BA FNP FPR F1 BA
Window_1 RF 15.83 0.01 91.30 92.08 72.83 0.03 42.37 63.57
GCN 16.88 2.25 76.15 90.44 40.22 2.18 50.69 78.80
($\downarrow$1.05) ($\downarrow$2.24) ($\downarrow$15.15) ($\downarrow$1.64) ($\uparrow$32.61) ($\downarrow$2.15) ($\uparrow$8.32) ($\uparrow$15.27)
Window_2 RF 12.08 0.01 93.47 93.95 69.51 0.03 46.30 65.23
GCN 14.79 0.64 87.30 92.29 57.32 0.13 57.38 71.27
($\downarrow$2.71) ($\downarrow$0.63) ($\downarrow$6.17) ($\downarrow$1.66) ($\uparrow$12.19) ($\downarrow$0.10) ($\uparrow$11.08) ($\uparrow$6.04)
Window_3 RF 12.29 0.15 92.32 93.78 68.89 0.44 42.75 65.34
GCN 17.92 1.53 79.68 90.28 64.44 0.30 48.85 67.63
($\downarrow$5.63) ($\downarrow$1.38) ($\downarrow$12.64) ($\downarrow$3.50) ($\uparrow$4.45) ($\uparrow$0.14) ($\uparrow$6.10) ($\uparrow$2.29)
Window_4 RF 11.25 0.37 91.22 94.19 51.38 0.00 65.43 74.31
GCN 18.96 2.54 73.40 89.25 42.20 0.40 69.23 78.70
($\downarrow$7.71) ($\downarrow$2.17) ($\downarrow$17.82) ($\downarrow$4.94) ($\uparrow$9.18) ($\downarrow$0.40) ($\uparrow$3.80) ($\uparrow$4.39)
Window_5 RF 13.54 0.86 86.46 92.80 64.89 0.61 48.94 67.25
GCN 15.62 3.26 71.68 90.56 39.69 0.72 70.85 79.79
($\downarrow$2.08) ($\downarrow$2.40) ($\downarrow$14.78) ($\downarrow$2.24) ($\uparrow$25.20) ($\downarrow$0.11) ($\uparrow$21.91) ($\uparrow$12.54)
Window_6 RF 22.71 0.17 85.88 88.56 84.93 0.21 25.14 57.43
GCN 22.50 2.45 71.81 87.53 43.15 0.66 66.14 78.09
($\downarrow$0.21) ($\downarrow$2.28) ($\downarrow$14.07) ($\downarrow$1.03) ($\uparrow$41.78) ($\downarrow$0.45) ($\uparrow$41.00) ($\uparrow$20.66)
Window_7 RF 9.79 3.03 75.90 93.59 16.58 0.25 89.91 91.59
GCN 17.92 2.57 73.85 89.76 18.98 1.45 83.70 89.78
($\downarrow$8.13) ($\uparrow$0.46) ($\downarrow$2.05) ($\downarrow$3.83) ($\downarrow$2.40) ($\downarrow$1.20) ($\downarrow$6.21) ($\downarrow$1.81)
Window_8 RF 15.00 1.10 84.04 91.95 84.95 0.15 25.23 57.45
GCN 19.79 1.76 77.23 89.23 54.84 0.49 56.76 72.34
($\downarrow$4.79) ($\downarrow$0.66) ($\downarrow$6.81) ($\downarrow$2.72) ($\uparrow$30.11) ($\downarrow$0.34) ($\uparrow$31.53) ($\uparrow$14.89)
Like in Section IV, we compare the performance measurements of the two models both before and after leakage apps are removed from testing data. Leakage portion is defined as $Leak_{\textit{fv}}\cup Leak_{\textit{emb}}$. Table III shows the results. For each testing window, the color-coded performance differences are included in parentheses, with green in favor of GCN and red in favor of RF. If we measure the performance on all the test data, RF consistently outperforms GCN based on the F1 score and balanced accuracy metric. However, once the leakage is removed, the comparative result flips: GCN outperforms RF in seven out of the eight testing windows, and by big margins. It is clear that after leakage is removed, not only both models’ performance decreases, but their comparative results also change to the opposite. In this case, the train-test leakage would have led to a completely different conclusion being drawn about the comparative advantage between the GCN and RF models on Android malware detection.
Therefore, train-test leakage can lead to two key issues: inflated research performance and misleading research conclusions. The inflated performance occurs because the model is inadvertently trained on information it will encounter during testing. This can skew the results and make the model appear more effective than it truly is. Consequently, these misleading results can influence research conclusions.
VI Which is Important: Memorization or Generalization?
In Section I we set out to answer two questions: 1) how will train-test leakage impact an ML model’s measured performance for Android malware detection, and 2) given that such leakage will happen “naturally” when the model is applied to real-world data, is there any practical value in discerning such differences of performance measurement? At this point we have clearly answered question 1). In this section we look at question 2). In our experiments we remove the leakage apps from testing data before evaluating a model’s performance. These leakage apps would be present in the real world data when the ML model is applied; wouldn’t evaluation on the complete test data (including the leakage apps) reflect the model’s utility in the real world? This question is especially relevant when research conclusion about different models’ comparative advantages flips when leakage data is removed (e.g., in case study 2). Should practitioners care about this result, if all they need is a model that would perform better on the real world data?
In real-world operation, all data must be processed. Removing the leakage from the test set is intended to evaluate the ML model’s capacity to generalize – to determine whether good performance is due to the model’s ability to generalize to unseen samples or simply because it has memorized the data it has already seen. This differentiation is important for deciding which ML model to use, or whether to use ML at all. It will be helpful to understand why the comparative results flipped in case study 2. Figure 5 illustrates the performance of RF and GCN models on 1) the complete test data, 2) the leakage portion of the test data, and 3) the non-leakage portion of the test data. Non-leakage portion is the test data after the leakage is removed. The figure shows that both models achieve their highest performance when tested on the leakage portion of the test data and the lowest on the non-leakage portion of test data. But their comparative differences are opposite between the two portions. RF outperforms GCN on the leakage portion, indicating that RF is better than GCN at memorization. GCN outperforms RF on non-leakage portion, indicating that GCN is better than RF at generalization. It turns out that RF’s advantage at memorization outweights its disadvantage on generalization, and shows stronger overall performance on the complete test data. The question is, does this mean RF should be recommended as having better utility for real-world operation? In other words, which capability is more important for the practical utility of a machine learning model: memorization or generalization? The answer is resoundinlgy the latter. This is because for apps that are known malware, there are other more efficient methods to detect them, such as through a known signature. Only when a malware has not been widely recognized a machine learning model could potentially provide useful guidance. Thus, evaluating the machine learning model on the non-leakage part of the testing data is more meaningful than evaluating on the complete testing data.
To further illustrate this point, we create a hypothetical detector, which we call “leak-aware detector,” for the purpose of approximating the effect of combining machine learning and signature-based approaches to detect Android malware. For an app the detector needs to make a decision on, we first determine if the app is considered leakage, i.e., is there an app in the training data that has identical or highly similar representation to it? If so we identify all such apps and use their labels’ majority vote to make a decision. Otherwise we use the trained machine learning model to make a prediction. This clearly separates generalization and memorization: we only use machine learning for generalizing, and use signature matching for memorizing. This leak-aware scheme more accurately reflects a machine learning model’s utility in practice. It is designed as an approximate model for combining machine learning with signature detection and used as a measurement tool only. It is by no means an indication on how such combination shall be implemented for operations. In fact it is not an efficient implementation since the detector needs to constantly keep track of the set of apps that have been used in training the machine learning model. We implemented such leak-aware detectors for the models in case study 2, to answer the question of “whether the flipped research conclusion makes any difference in the real world.” For each model – RF and GCN, we utilize the same leakage definition – $Leak_{\textit{fv}}\cup Leak_{\textit{emb}}$ to build the leak-aware detector described above.
Figure 5: Performance comparison of RF and GCN on the leakage and non-leakage portions of the test data, as well as on the complete test data.
Figure 6: Leak-aware vs. Stand-alone Detector Performance
The performance results of the two leak-aware detectors, along with the stand-alone versions (applying the ML model on all test data), are shown in Figure 6. The leak-aware GCN detector’s overall performance on the complete test data exceeds both the stand-alone RF and the leak-aware RF. This is in contrast to the comparison of stand-alone GCN and stand-alone RF, where RF outperforms GCN on the complete test data.
This result shows that not only GCN outperforms RF on the testing data with leakage removed, it also outperforms RF on the complete test data when used in the leak-aware scheme. Thus, the flipped comparative result between the two models is not only relevant to understanding the models’ ability to generalize to data it has not seen, it also has practical implications when the models are applied to real-world data and in combination with signature-based detection schemes. Specifically, by approximating the effect of such combination, the leak-aware detector allows the GCN’s comparative strength against RF to show on the complete testing data as well.
VII Related Work
Kapoor and Narayanan [1] systematically examine the leakage problem in machine learning-based science. The authors provide a taxonomy of data leakage with eight categories. The train-test leakage discussed in this paper belongs to the category [L1.4] Duplicates in datasets. The authors also conducted a comprehensive survey of literature and found 17 fields, computer security included, where researchers have published papers revealing data leakage problems in machine learning-based science. The security paper cited by Kapoor and Narayanan is Arp et. al 2022 [10], which covers all the other seven types of leakage except [L1.4]. We have also extensively searched the computer security literature. The only mentioning of train-test leakage we could find is from a recent manuscript by Ding et. al [24], in the context of vulnerability datasets for evaluating code language models. The authors identify both train-test leakage (called “Code Copy”) and temporal leakage (called “Time Travel”) in those datasets. The paper shows the percentage of train-test leakage in these datasets, but did not conduct experiments to specifically analyze these leakages’ impact on evaluating code language models’ performance. Allamanis [30] analyzed the adverse effects of code duplication in machine learning models of code. While not explicitly using the term “data leakage,” the “cross-set duplicates” described in the paper is essentially train-test leakage. There are also works in other domains as discussed earlier [2, 3, 4] that explicitly use the term “data leakage” to describe this phenomenon.
Monlina et al. [31] pointed out the redundancy problem in datasets used for machine learning based Android malware detection. The authors conduct experiments on the impact of duplicate samples in both training and testing data sets, but did not analyze the impact of duplicates between training and testing set, which is the train-test leakage we focus on. The authors through experiments point out that duplicates in training data result in biased model, whereas duplicates in testing data result in biased evaluation result that disproportionately emphasizes prediction results on those duplicate samples. The authors used an approach proposed by Surendran [32] to filter the dataset – removing duplicates and retaining only one representative for each duplicate group, so that the trained model can generalize better and the measured performance can better indicate generalization ability of detectors. Our work examines the data leakage between training and testing sets, which is an orthogonal problem to duplicates within training and testing sets. Our results thus complement those from this prior work.
The train-test leakage discussed in this paper arises from the presence of identical or nearly identical app representations for distinct apps. This strongly indicates that the raw APK files share significant similarities. Extensive studies have previously delved into Android app similarities [33, 34, 35, 36, 37, 38, 39, 40, 41], with the primary focus of similarity measurement on localizing and detecting changes resulting from app repackaging [33], targeting app clones or code clones [36], identifying changes among app versions to verify valid code updates [35], and detecting vulnerable code reuse in malware variants [34].
VIII Conclusion
This work highlights a prevalent issue of train-test data leakage in machine learning (ML)-based Android malware detection. We demonstrate this issue through two case studies, illustrating that such leakage can inflate model performance and sometimes also lead to misleading research conclusions. The key issue is that the existence of train-test leakage causes the evaluation results to conflate a model’s ability to generalize with its ability to memorize. Thus it is more meaningful to evaluate a machine learning model’s performance on testing data with the leakage removed, than on the complete testing data. We further use a leak-aware detector scheme to approximate the effect of combining machine learning with signature-based detection, and shows more clearly which model can provide more practical utility in real-world use scenarios.
Appendix A: Evaluation Metrics Definition
First some common terminologies:
TP: True Positives (correctly classified malicious apps)
FN: False Negatives (malware mis-classified as benign)
TN: True Negatives (correctly classified benign apps)
FP: False Positives (benign apps mis-classified as malware)
False Positive Rate (FPR) and False Negative Rate (FNR) quantify the rate at which the the model incorrectly classify a sample:
$\textit{FPR}=\frac{\textit{FP}}{\textit{TN}+\textit{FP}}\hskip 14.45377pt\textit{FNR}=\frac{\textit{FN}}{\textit{TP}+\textit{FN}}$
FNR and FPR are insensitive to the class ratio between malicious and benign samples. In practice the small ratio of malware relative to benign apps makes these metrics unsuitable to reflect a model’s real utility, due to the famous base-rate fallacy [42]. For such imbalanced data, precision and recall, and their combination F1 score, are more suitable.
$\textit{Precision}=\frac{\textit{TP}}{\textit{TP}+\textit{FP}}\hskip 14.45377pt\textit{Recall}=\frac{\textit{TP}}{\textit{TP}+\textit{FN}}$
[…]
Appendix B: Details about Our GCN Model
Our GCN model utilizes a four-layer graph convolutional architecture followed by a standard linear layer from the torch.nn library. ReLU (Rectified Linear Unit) activation ($max(x,0)$) was employed in the convolutional layers to extract localized node embeddings. We experimented with varying the number of convolutional layers (2-7) and hidden channels (32-128) within the model. While increasing hidden channels improved efficiency, GPU memory limitations prevented exploration beyond that. We found optimal performance with four convolutional layers and 64 hidden channels. Hyperparameter tuning identified 200 epochs and a learning rate of 0.001 as optimal for our dataset.
Appendix C: Embedding Similarity Threshold
In Section V, $M$ defines the level of similarity between graph embeddings for the purpose of considering train-test leakage. The following details the process of establishing this threshold for each data window.
For a given data window, an $M$ value will give rise to a subset of apps $\mathbf{Leak_{\textit{emb}}}$ in the testing data that are considered leakage for the GCN model. On the other hand, the subset $\mathbf{Leak_{\textit{fv}}}$ captures the leakage for the RF model. Intuitively, being leakage means that a testing app is very similar to a training app, whether under the RF model or the GCN model. Thus we’d like to choose an $M$ so that the corresponding subset $\mathbf{Leak_{\textit{emb}}}$ most closely matches the subset $\mathbf{Leak_{\textit{fv}}}$. We use the metric Intersection over Union (IoU) to measure how closely the two sets are aligned:
$\textit{Intersection over Union}\left(\textit{IoU}\right)=\frac{\mathbf{Leak_{\textit{fv}}}\cap\mathbf{Leak_{\textit{emb}}}}{\mathbf{Leak_{\textit{fv}}}\cup\mathbf{Leak_{\textit{emb}}}}$
IoU represents the proportion of elements that are common to $\mathbf{Leak_{\textit{fv}}}$ and $\mathbf{Leak_{\textit{emb}}}$ relative to their total combined size. It is a normalized score between 0 and 1. When the two sets completely coincide, IoU is 111. When the two sets are disjoint, IoU is 0.
We follow the steps below to find the threshold $M$ that maximizes IoU:
  1. 1.
    Define a similarity range (e.g., [0.8, 1.0]).
  2. 2.
    Set a small increment rate (e.g., 0.01) to step through the range.
  3. 3.
    At each step, calculate the IoU based on the current threshold value $M$.
  4. 4.
    We identify the $M$ that leads to the maximum IoU.
The above steps can be achieved through Algorithm 1. Figure 7 illustrates how we find thethreshold $M$ for a given experiment window.
Algorithm 1 Finding the Optimal Threshold $M$
1:Define a similarity range $R$ (e.g., [0.8, 1.0]).
2:Set a small increment rate r (e.g., 0.01) to step through the range.
3:Initialize $M_{\text{max}}$ to 0, $\text{IoU}_{\text{max}}$ to 0
4:for $M$ in range from $R_{\text{min}}$ to $R_{\text{max}}$ with step $r$ do
5:  Calculate IoU from current $M$ and $Leak_{\textit{emb}}$.
6:  if $\text{IoU}>\text{IoU}_{\text{max}}$ then
7:   $\text{IoU}_{\text{max}}$ = IoU
8:   $M_{\text{max}}$ = $M$
9:  end if
10:end for
11:Output: The optimal threshold $M_{\text{max}}$ that leads to the maximum IoU.
Figure 7: Determining the cosine similarity threshold ($M$) within the experiment window. $M$ is the value that leads to the maximum IoU.
Win_1 Win_2 Win_3 Win_4 Win_5 Win_6 Win_7 Win_8
Default Test 8,000 8,000 8,000 8,000 8,000 8,000 8,000 8,000
$\mathbf{Leak_{\textit{fv}}}$ 3,245 3,001 3,134 3,865 4,297 3,030 2,132 2,552
$\mathbf{Leak_{\textit{emb}}}$ 4,141 3,585 4,181 4,701 5,514 3,180 3,271 4,364
$\mathbf{Leak_{\textit{fv}}\cup Leak_{\textit{emb}}}$ 4,694 4,203 4,929 5,362 6,066 4,529 4,395 5,256
TABLE IV: Statistics of test data on each window. $\mathbf{Leak_{\textit{emb}}}$ is determined for each experiment window by selecting the $M$ value that maximizes the IoU score.
Train Test Win_1 Win_2 Win_3 Win_4 Win_5 Win_6 Win_7 Win_8
Win_1 0.4056 0.2699 0.1719 0.1596 0.1535 0.1301 0.0568 0.0321
Win_2 0.3751 0.2256 0.2116 0.1900 0.1585 0.0636 0.0405
Win_3 0.3917 0.3275 0.2670 0.1874 0.0811 0.0549
Win_4 0.4831 0.3698 0.2285 0.1140 0.0620
Win_5 0.5371 0.2850 0.1295 0.0770
Win_6 0.3787 0.1565 0.1003
Win_7 0.2665 0.1224
Win_8 0.3190
TABLE V: $Leak_{\textit{fv}}$ ratio in different test window while keep the training data fixed. A decreasing ratio of $Leak_{\textit{fv}}$ as the test data release timeline progresses.
Table IV summarizes the appearance of $\textit{Leak}_{\textit{fv}}$ or $\textit{Leak}_{\textit{emb}}$ in our test data. Table V shows a consistent decline in the $\textit{Leak}_{\textit{fv}}$ ratio within the test data over time, considering a fixed training dataset. For instance, training models solely on window-1 data leads to progressively lower $\textit{Leak}_{\textit{fv}}$ occurrences in newer test datasets. The $\textit{Leak}_{\textit{emb}}$ ratio follows a similar trend.
Appendix D: Why So Much Leakage?
Multiple factors could contribute to identical or nearly identical Android app representations. One is the rapid iteration of app versions. The other reason is the region-based app release: one app could have multiple language versions for different regions. While many APK files are distinct entities based on hash values, they often represent the same application with different versions. Another factor is the reuse of software components during the software development process. Additionally, the method or data selected to represent the app is another contributing factor.
It is surprising though to see such high percentage of testing apps that share the same feature vectors with training apps (Figure 2 and 4), especially since these models use thousands and even tens of thousands of features. To delve deeper into the phenomenon of identical feature vectors, we manually analyzed four diverse sets of apps. Each set comprised at least one malicious and one benign app, sharing an identical feature vector within the set. We randomly selected these sets to ensure they encompass a representative range of potential scenarios:
[…]
e9060bf176250385b264777dc49ec5fea6a8df7ce2d52fc103553fca0b5c6f41.apk,
[…]
293652541d80f4b9ecb0b60a9cd89abeff5c7d04d909bcd04cefef7ba6e25478.apk,
[…]
4cfbe089ab61c4254308375e9a3fcf13eaa52f22d7810aecaf09ab7b66a6984d.apk,
[…]
9215b52ed735b38d4fb91fe3f12af4d2d18f596752a028aeebaf4049e8dca464.apk,
[…]
The identical feature vectors observed in sets $S_{1}$, $S_{2}$, and $S_{4}$ can be attributed to a common factor: all the apps in these sets were developed using the Seattlecloud app development framework (https://sc.mobileappframeworks.com/). This framework apparently leads to significant similarities in manifest files and APIs across apps, even those with diverse functionalities, such as a paperfolding-crafts app and a Spanish radio station (examples from set 1). Our investigation revealed that the primary difference between these apps built using the same platform lies in their HTML code, which is currently not considered by our feature vector. Notably, some of these apps utilize inline frames (iframes) within their HTML, which trigger malicious flags from certain security vendors. Further investigation is needed to determine whether these flags are false positives.
Following a similar pattern to the previous three sets, all apps in set $S_{3}$ are built using the Cordova development framework. Despite serving distinct purposes (e.g., one app for Arabic scripts/preaching and another for Perdana 4D, a lottery company), these apps exhibit remarkably similar manifest files and API calls. Once again, the key differentiating factor lies in their HTML code. Interestingly, some apps with suspicious URLs within their HTML trigger “malicious” flags from certain anti-malware vendors. Again, further investigation is needed to determine if these flags are false positives.
Our analysis reveals a growing trend: developers are increasingly utilizing well-established app development frameworks to enhance development efficiency. This trend, however, has an unintended consequence: apps with diverse functionalities can end up exhibiting remarkably similar manifest files, API calls, and even bytecode. The key functionalities of these apps are primarily implemented through their HTML code, while the bytecode serves more as a common framework for the HTML code to execute its specific functions within each app.
This section only reports our anectdotal findings on the causes of the significant train-test leakage in Android app datasets. More thorough study is needed to fully understand all causes and is beyond the scope of this paper. Our results do call for awareness of the train-test leakage phenomenon in this problem domain, and necessary steps be taken to remove the leakage’s impact on performance evaluation. Moreover, it is perhaps time for a more careful consideration on how to represent Android apps for machine learning based malware detection, given that apps with distinct functionalities could end up with very similar representations under the scheme commonly used by the current research methods.